# Restricted Spectral Gap Decomposition for Simulated Tempering Targeting Mixture Distributions

### v1

@@ C.2.3. Auxiliary Lemmas
[…]
Hence, the probability of transitioning from state $(i,j)$ to $(i-1,j)$ under the projected chain $\widehat{M}$ is given by
@@
    Let $z=(i,j)$ and $w=(i-1,j)$. Then the edge $(z,w)$ is used only by paths between vertices $x$ and $y$ such that one lies in the set $S:=\{i,\ldots,L\}\times\{j\},\text{and the other in }S^{c}.$ Therefore, its contribution to the edge congestion is bounded by
@@
    Let $z=(1,j)$ and $w=(1,j^{\prime})$. Then the edge $(z,w)$ is used only by paths between vertices $x$ and $y$ such that one of them lies in the set $[L]\times\{j\}$ and the other in $[L]\times\{j^{\prime}\}$. Therefore, its contribution to edge congestion is bounded by
@@ C.2.5. Restricted Spectral Gap Bound
We now invoke Theorem 2 to bound the $[L]\times\mathcal{X}^{0}$-restricted spectral gap of the approximate STMH chain $\widetilde{M}$, as formalized in the next lemma.
@@
Let $\widetilde{M}$ denote the approximate STMH chain, as defined in Definition 5, with $\lambda$ being a fixed constant. Under the same conditions as in Lemma 19, the $[L]\times\mathcal{X}^{0}$-restricted spectral gap of $\widetilde{M}$ admits the following lower bound
@@ C.3. Restricted Spectral Gap of the STMH chain
We now establish a lower bound on the $[L]\times\mathcal{X}^{0}$-restricted spectral gap of the STMH chain $M^{*},$ as formalized in next lemma.
@@
Let $M^{*}$ be the STMH chain, as defined in Definition 4, with $\lambda$ being a fixed constant. Under the same conditions as in Lemma 19, the $[L]\times\mathcal{X}^{0}$-restricted spectral gap of $M^{*}$ admits the lower bound
@@ C.4.1. Assumptions on the Parameters
We now describe how to choose the algorithm parameters—number of temperature levels $L$, inverse temperature sequence $(\beta_{i})_{i=1}^{L}$, temperature-swap rate $\lambda$, proposal step size $\eta$, initial covariance matrix $\Sigma_{0}$, number of iterations $N$—so that the STMH algorithm achieves the asserted time complexity. Recall that $\kappa=\gamma_{\max}/\gamma_{\min}$.
@@
[…]
@@ C.4.2. Auxiliary Lemmas
Taking logarithms and using $\log(1+x)\leq x$ for all $x>-1$, we obtain
[…]
@@
From Equation (C.25), we have fixed constants $20<s_{1},s_{2}<$ 2 such that
@@
Under the assumptions of the lemma, and by Lemmas 21 and 22, the []×$[\ell]\times\mathcal{X}^{0}$-restricted spectral gap of $M^{*}$ satisfies
@@
where $C^{\prime},c^{\prime}>0$ are fixed constants. Next, by applying Lemma 29 with constants $C={1}/{2e}$ and $c={1}/{L}$, we obtain the following bound
[…]
@@ C.4.3. Proof of Theorem 5
Let $L$ denote the number of temperature levels defined in Equation (C.20). By applying Lemma 28 inductively with $\delta=\varepsilon/(4L)$, we obtain that, with probability at least $41-\varepsilon/$ 4, the following bound holds
@@
To ensure this guarantee, it suffices to generate $s=L^{2}\log\left({4L}/{\varepsilon}\right)$ samples from each temperature level $i\in[L]$, resulting in a total of $sL=L^{3}\log\left({4L}/{\varepsilon}\right)$ samples from Algorithm 1. Applying Lemma 27 with $\alpha={\varepsilon}/{(4L^{4}\log(4L/\varepsilon))}$, we obtain that, with probability at least $41-\varepsilon/$ 4, we obtain $s$ samples from each temperature level $i\in[L]$ by running Algorithm 1 for $N$ steps (as defined in Equation (C.24)) and repeating this process independently $T$ times, where
@@
where $c^{\prime},C^{\prime}>0$ are fixed constants. By applying Lemma 26 and Lemma 27, we conclude that, with probability at least $41-\varepsilon/$ 4, Algorithm 1 produces a sample from a distribution that is within total variation distance $\varepsilon/4$ of the target distribution $P^{*}$ in time $T_{\text{sampling}}$, where

### v2

@@ C.2.3. Auxiliary Lemmas
$\widetilde{a}\big((i,j,x),(i^{\prime},j,x)\big)=\min\left\{\frac{r_{i^{\prime}}\,\widetilde{p}_{(i^{\prime},j)}(x)}{r_{i}\,\widetilde{p}_{(i,j)}(x)},\;1\right\}.$
Hence, the probability of transitioning from state $(i,j)$ to (i1,j)(i-1,j) under the projected chain $\widehat{M}$ is given by
@@
    Let $z=(i,j)$ and w=(i1,j)w=(i-1,j). Then the edge $(z,w)$ is used only by paths between vertices $x$ and $y$ such that one lies in the set $S:=\{i,\ldots,L\}\times\{j\},\text{and the other in }S^{c}.$ Therefore, its contribution to the edge congestion is bounded by
@@
    Let $z=(1,j)$ and $w=(1,j^{\prime})$. Then the edge $(z,w)$ is used only by paths between vertices $x$ and $y$ such that one of them lies in the set $L]\times\{j\}$ and the other in $L]\times\{j^{\prime}\}$. Therefore, its contribution to edge congestion is bounded by
@@ C.2.5. Restricted Spectral Gap Bound
We now invoke Theorem 2 to bound the $L]\times\mathcal{X}^{0}-$restricted spectral gap of the approximate STMH chain $\widetilde{M}$, as formalized in the next lemma.
@@
Let $\widetilde{M}$ denote the approximate STMH chain, as defined in Definition 5, with $\lambda$ being a fixed constant. Under the same conditions as in Lemma 19, the $L]\times\mathcal{X}^{0}-$restricted spectral gap of $\widetilde{M}$ admits the following lower bound
@@ C.3. Restricted Spectral Gap of the STMH chain
We now establish a lower bound on the $L]\times\mathcal{X}^{0}-$restricted spectral gap of the STMH chain $M^{*},$ as formalized in next lemma.
@@
Let $M^{*}$ be the STMH chain, as defined in Definition 4, with $\lambda$ being a fixed constant. Under the same conditions as in Lemma 19, the $L]\times\mathcal{X}^{0}-$restricted spectral gap of $M^{*}$ admits the lower bound
@@ C.4.1. Assumptions on the Parameters
We now describe how to choose the algorithm parameters—number of temperature levels $L$, inverse temperature sequence $(\beta_{i})_{i=1}^{L}$, temperature-swap rate $\lambda$, proposal step size $\eta$, initial covariance matrix $\Sigma_{0}$, number of iterations NN—so that the STMH algorithm achieves the asserted time complexity. Recall that $\kappa=\gamma_{\max}/\gamma_{\min}$.
@@
(C.22)  $\quad\quad\quad\text{ where }\nu=1+\log\kappa+\frac{2}{d}\log\left(\frac{2}{w_{\min}}\right),$
@@ C.4.2. Auxiliary Lemmas
Taking logarithms and using $\log(1+x)\leq x$ for all x>1x>-1, we obtain
$\log C_{L}=(L-1)\,\log\,\!\Bigl(1+\tfrac{2}{\,L-1}\Bigr)\;\leq\;(L-1)\Bigl(\tfrac{2}{\,L-1}\Bigr)\;\leq\;2,$
@@
From Equation (C.25), we have fixed constants $0<s_{1},s_{2}<2$ such that
@@
Under the assumptions of the lemma, and by Lemmas 21 and $2$, the $\ell]\times\mathcal{X}^{0}-$restricted spectral gap of $M^{*}$ satisfies
@@
In the following lemma, we analyze how many times Algorithm 1 must be re-run, with a fixed number of steps $N$, in order to obtain a sample from the desired temperature level.
@@
where $C^{\prime},c^{\prime}>0$ are fixed constants. Next, by applying Lemma 29 with constants C=1/2eC={1}/{2e} and c=1/Lc={1}/{L}, we obtain the following bound
$\left|\frac{\widehat{Z}_{\ell+1}/Z_{\ell+1}}{\widehat{Z}_{\ell}/Z_{\ell}}\right|\in\left[1-\frac{1}{L},\;1+\frac{1}{L}\right].$
@@ C.4.3. Proof of Theorem 5
Let $L$ denote the number of temperature levels defined in Equation (C.20). By applying Lemma 28 inductively with $\delta=\varepsilon/(4L)$, we obtain that, with probability at least $1-\varepsilon/4$, the following bound holds
@@
To ensure this guarantee, it suffices to generate $s=L^{2}\log\left({4L}/{\varepsilon}\right)$ samples from each temperature level $i\in[L]$, resulting in a total of $sL=L^{3}\log\left({4L}/{\varepsilon}\right)$ samples from Algorithm 1. Applying Lemma 27 with $\alpha={\varepsilon}/{(4L^{4}\log(4L/\varepsilon))}$, we obtain that, with probability at least $1-\varepsilon/4$, we obtain $s$ samples from each temperature level $i\in[L]$ by running Algorithm 1 for $N$ steps (as defined in Equation (C.24)) and repeating this process independently $T$ times, where
@@
where $c^{\prime},C^{\prime}>0$ are fixed constants. By applying Lemma 26 and Lemma 27, we conclude that, with probability at least $1-\varepsilon/4$, Algorithm 1 produces a sample from a distribution that is within total variation distance $\varepsilon/4$ of the target distribution $P^{*}$ in time $T_{\text{sampling}}$, where